a trapezoid with area 152 has a longer base of length 22. The shorter base and the height have a ration1:2. Find the height of the trapezoid
Answer by Boreal(15235)

A=(1/2)h(b1+b2); b1=x, b2=22
152=(1/2)2x(x+22)
304=2x^2+44x
2x^2+44x-304=0
x^2+22x-152=0
x=(1/2)(-22+/- sqrt (484+608); sqrt (1092)=33.05
positive root only.
x=5.523
bases are 5.523, 22, and height is 11.046
area is (1/2)*11.046*(27.523)=152.01 (rounding error)
height is 11.046